Wescott Company has three divisions: A, B, and C. The company has a hurdle rate of 8 percent. Selected operating data for the three divisions are as follows:

Division A Division B Division C
Sales revenue $ 1,215,000 $ 1,262,000 $ 1,294,000
Cost of goods sold 751,000 927,000 940,000
Miscellaneous operating expenses 82,000 70,000 71,000
Interest and taxes 66,000 59,000 59,000
Average invested assets 11,414,000 2,776,000 4,563,000


Wescott is considering an expansion project in the upcoming year that will cost $7.1 million and return $642,000 per year. The project would be implemented by only one of the three divisions.


Required:
1.
Compute the ROI for each division. (Do not round your intermediate calculations. Round your percentage answer to 2 decimal places, (i.e. 0.1234 should be entered as 12.34%.))

ROI A_______

ROI B_______
ROI C_______





2. Compute the residual income for each division. (Loss amounts should be indicated by a minus sign.)

Residual income

Division A_____

Division B______

Division C_______





3. Rank the divisions according to the ROI and residual income of each.

Rank

Division A_____

Division B______

Division C______





4-a. Compute the return on investment on the proposed expansion project. (Round your percentage answer to 2 decimal places, (i.e. 0.1234 should be entered as 12.34%.))

Return on investment of proposed expansion project____%



4-b. Is this an acceptable project?

No
Yes



5. Without any additional calculations, state whether the proposed project would increase or decrease each division’s ROI.

Division A_____

Division B______

Division C_______

6. Compute the new ROI and residual income for each division if the project was implemented within that division. (Loss amounts should be entered with a minus sign. Enter your ROI percentage answers to 2 decimal places, (i.e., 0.1234 should be entered as 12.34%.))

ROI %

A
B
C

Residual Income (loss)

A
B
C
Answer #1

ROI = Net income / average invested assets

ROI of Division A = 316000/11414000 * 100

= 2.77%

ROI of Division B = 206000 / 2776000 * 100

= 7.42%

ROI of Division C = 224000/4560000 * 100

= 4.91%

Working:

1. Net income of each division = sales - cost of goods sold - expenses

Division A = $1215000 - 751000 -82000 - 66000

= $316000

Division B = $1262000 - 927000 - 70000 -59000

= $206000

Division C = $1294000 - 940000 - 71000 - 59000

= $224000

2. Residual income = Net operating income - (required rate * average assets)

Net operating income = Net income + Interest and taxes

Division A = ($316000+66000) - 8%*11414000

= $382000 - 913120

= -$531120

Division B = ($206000 + 59000) - 8%*2776000

= $265000 - 222080

= $42920

Division C = ($224000 +59000) - 8%*4563000

= -$64300 = $283000 - 365040

= -$82040

3. Rank as per ROI:

Rank 1 division B

Rank 2 division C

Rank 3 Division A

Rank as per Residual income:

Rank 1 Division B

Rank 2 division C

Rank 3 Division A

4. Return on proposed project = 642000/7100000 * 100 = 9.04%

4b Yes

5Due to ROI of the new project is high as compared to earlier, that implies the proposed project increase each decision ROI.

6. ROI of Division A = (316000+642000) /(11414000 +7100000) * 100

= 958000/18514000*100 = 5.45%

ROI of Division B = (206000+642000) /(2776000+7100000) * 100

= 848000/ 9876000*100 =8.59%

ROI of Division C =( 224000+642000) /(4563000 + 7100000) * 100

= 866000/11663000*100 = 7.43%

Residual income = Net operating income - (required rate * average assets)

Net operating income = Net income + Interest and taxes

Division A = $958000 - 8%*18514000 = $958000 - $1481120

= -$523120

Division B = $848000 - 8%*9876000 = $848000 - 790080

= $57920

Division C = $866000 - 8%*11663000 = $866000 - $933040

= -$67040